OpenStudy (sstradic):

A cone-shaped paper cup can hold 35 cubic inches of food. If the radius of the cone is 2.4 inches, what is the height of the cone? Use 3.14 for pi. Round your answer to the nearest hundredth. 5.81 inches 6.57 inches 7.89 inches 8.89 inches

OpenStudy (sstradic):

plz help

OpenStudy (campbell_st):

so the formula for volume of a cone is \[V = \frac{1}{3}\pi r^2 h\] you know V = 35, r = 2.4 and pi = 3.14. so substitute the values and solve for h